OpenStudy (anonymous):

Solve the system of equations. { x+6y=-23 {9x+5y=-60

OpenStudy (anonymous):

x=-23-6y Plug in this value for x in the other equation. 9(-23-y)+5y=-60 Simplify. -207-9y+5y=-60 -207-4y=-60 -4y=147 y=-36.75 Substitute y back into the first equation. x=-23-6(-36.75) x=-23+220.5 x=197.5

OpenStudy (anonymous):

So is it (-6,-2) (5,-2) (-5,-3) Undefined

OpenStudy (anonymous):

ktruman, you need to look at your substitution again. It should be 23-6y, not -23-y.